OpenStudy (anonymous):

An object is dropped from a height of 25 meters. At what velocity will it hit the ground? A. 7.0 meters/second B. 11 meters/second C. 22 meters/second D. 49 meters/second E. 70 meters/second

OpenStudy (michele_laino):

here we have to apply the formula of Torricelli, namely the requested velocity v, is given by the subsequent formula: \[v = \sqrt {2 \times g \times h} \] where g=9.81 m/sec^2

OpenStudy (michele_laino):

\[v = \sqrt {2 \times g \times h} = \sqrt {2 \times 9.81 \times 25} = ...?\]

OpenStudy (anonymous):

so sqrt(2 X 9.8 X 25?

OpenStudy (michele_laino):

yes!

OpenStudy (anonymous):

so 22.13 got it

OpenStudy (michele_laino):

that's right! Now you have to round off that value, so what do you get?

OpenStudy (anonymous):

22 thanks alot

OpenStudy (michele_laino):

thanks!

OpenStudy (anonymous):

no seriously thanks this was giving me trouble for a while and all the other places i checked gave me an incomplete equation to work with

OpenStudy (michele_laino):

ok! :)